OpenStudy (anonymous):

Find the exact value of the expression. If tan θ = 9, find the exact value of cot ((pi/2)-θ) A. 8 B. 10 C. 1/9 D. 9

OpenStudy (anonymous):

use tan(x)=sin(x)/cos(x); cot(x)=cos(x)/sin(x); cos (pi/2-x)=sin(x) ; sin(pi/2-x)=cos(x)

OpenStudy (anonymous):

This is a double indirection kind of thing. tan(x) = 1/cot(x) tan(pi/2-x) = 1/tan(x)

OpenStudy (anonymous):

So the inverse of the inverse is ?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

@telliott99

OpenStudy (anonymous):